Find (f*g)(x) and (g*f)(x) for each f(x) and g(x)
f(x)=x+5
g(x)=x-3

Answer by stanbon(75887)

Find (fog)(x) and (gof)(x) for each f(x) and g(x)
f(x)=x+5
g(x)=x-3
----------
fog(x) = f[g(x)] = f[x-3] = (x-3+5) = x+2
-------------
gof(x) = g[f(x)] = g[x+5] = x+5-3 = x+2
====================
Cheers,
Stan H.
====================